Birth weights at a local hospital have a normal distribution with a mean of 110 oz. and a standard deviation of 15 oz. What is the probability that a child born at this hospital weighs less than 7 lbs?
Answer by stanbon(75887)

Birth weights at a local hospital have a normal distribution with a mean of 110 oz. and a standard deviation of 15 oz. What is the probability that a child born at this hospital weighs less than 7 lbs?
------------------------------
7 lb = 7*16 = 112 oz
---
z(112) = (112-110)/15 = 2/15
-----------------
P(x < 7 lbs) = P(x < 112 oz)
= P(z < 2/15)
= normalcdf(-100,2/15)
= 0.5530
===========================
Cheers,
Stan H.
=============